# Reliability and Interpretability in Science and Deep Learning

### v2

@@ Appendix A On the irreducible complexity of DNN representations
One may question the meaningfulness of assessing the complexity of the assumptions by counting the words used to express them. In fact, the length of the expression of a set of axioms is language-dependent. Even worse, for any set of axioms, one can always find a language where the same axioms assume a trivially concise form (Kelly, 2009; Votsis, 2016). This is the origin of an old paradox: why do scientists employ relatively complex formulations, even for their most fundamental theories, when a much simpler one is available? Is it due to hidden cultural bias? An answer was given in Scorzato (2013), where it was shown that the language that makes the axioms trivially short typically employs concepts that are not measurable. If we require the axioms to include sufficient relations to measurable concepts, finding a concise formulation becomes non-trivial and the shortest formulation among all available languages becomes meaningful. In fact, the standard scientific formulations are typically the most concise available, under these constraints [^25] Indeed, the scientific language evolves with science in a global optimisation process (Kvasz, 2008)..
@@
Note that we have not proved that any DNN model cannot be represented with coordinates that are both very concise and also measurable. We have only shown that an easy, very concise, formulation is not measurable. But this is sufficient for our goals, namely to refute the claim that there is always an easy way to formally trivialise the formulations of any system of axioms (Kelly, 2009; Votsis, 2016). Although that claim remains correct from a logical point of view, it does not hold anymore once the constraints of measurability are enforced. Discovering a simpler formulation might be possible and it would be a significant achievement [^26] Note also that, in general, it is not possible to tell whether a given formulation is the shortest possible: this question is not computable, as shown in Chaitin (1969). However, what matters to judge the complexity of a real scientific model is the shortest formulation available, which certainly exists and it is what Def. 1 refers to..
@@ Appendix B A simple model of progress
A natural way to quantify the simplicity of the assumptions of a model is via Kolmogorov complexity (KC) (Li and Vitányi, 2019; Kolmogorov, 1965; Chaitin, 1969; Solomonoff, 1964; Hutter, 2005; Votsis, 2016). In general, the KC of a data string $D$ is defined as the length of the shortest program $p$ (in a given programming language) that produces $D$. In this context, the KC simply measures the length of the assumptions of the model ${\cal M}$ in the shortest formulation available in the given formal language (Chaitin, 1975; Scorzato, 2013). Comments about other frameworks can be found in the references cited in the App. C.
@@
The (epistemic) complexity of a model ${\cal M}$ is the minimal length—across all available formulations (in any language) of ${\cal M}$—of all the assumptions of ${\cal M}$ that are needed to derive all the available results of ${\cal M}$. The assumptions must contain all references needed to ensure that any comparison between the empirical data $D$ and the corresponding results of ${\cal M}$ are measurable [^30] For a discussion of the concept of “measurable” see App. A and (Scorzato, 2013).. The (epistemic) simplicity of a model ${\cal M}$ is the inverse of its complexity.
Being defined as the minimum across all available formulations, Def. 1 is obviously language independent. The interesting observation is that it is also in non-trivial, in general, thanks to the analysis in App. A.
@@ Appendix C Alternative notions of complexity
Kolmogorov complexity is a very popular measure of complexity, also widely used to estimate the complexity of a theory (Chaitin, 1975). Nevertheless, it is not the only possibility and one may question how the present analysis would change if we had used a different measure. For example, any specific compression software, like, e.g., gzip, implicitly defines an alternative notion of complexity. However, in all these cases, our conclusions about the complexity of DNN models would not change. Counting the atomic sentences (Carey, 2023) would also not change our conclusions.

### v3

@@ Appendix A On the irreducible complexity of DNN representations
One may question the meaningfulness of assessing the complexity of the assumptions by counting the words used to express them. In fact, the length of the expression of a set of axioms is language-dependent. Even worse, for any set of axioms, one can always find a language where the same axioms assume a trivially concise form (Kelly, 2009; Votsis, 2016). This is the origin of an old paradox: why do scientists employ relatively complex formulations, even for their most fundamental theories, when a much simpler one is available? Is it due to hidden cultural bias? An answer was given in Scorzato (2013), where it was shown that the language that makes the axioms trivially short typically employs concepts that are not measurable. If we require the axioms to include sufficient relations to measurable concepts, finding a concise formulation becomes non-trivial and the shortest formulation among all possible languages becomes meaningful. In fact, the standard scientific formulations are typically the most concise that we can find, under these constraints [^30] Indeed, the scientific language evolves with science in a global optimisation process (Kvasz, 2008)..
@@
Note that we have not proved that any DNN model cannot be represented with coordinates that are both very concise and also measurable. We have only shown that an easy, very concise, formulation is not measurable. But this is sufficient for our goals, namely to refute the claim that there is always a way to formally trivialise the formulations of any system of axioms (Kelly, 2009; Votsis, 2016). Although that claim remains correct from a logical point of view, it does not hold anymore once the constraints of measurability are enforced. Discovering a simpler formulation might be possible and it would be a significant achievement [^31] Note also that, in general, it is not possible to tell whether a given formulation is the shortest possible: this question is not computable, as shown in Chaitin (1969). However, that simply means that the complexity of a real scientific model can be only estimated approximatively, as it is the case for most scientific quantities..
@@ Appendix B A simple model of progress
A natural way to quantify the simplicity of the assumptions of a model is via Kolmogorov complexity (KC) (Kolmogorov, 1965; Chaitin, 1969; Solomonoff, 1964; Zenil, 2020a; Hutter, 2005; Votsis, 2016; Li and Vitányi, 2019). In general, the KC of a data string $D$ is defined as the length of the shortest program $p$ (in a given programming language) that produces $D$. In this context, the KC simply measures the length of the assumptions of the model ${\cal M}$ in the shortest formulation available in the given formal language (Chaitin, 1975; Scorzato, 2013). Comments about other frameworks can be found in the references cited in the App. C.
@@
The (epistemic) complexity of a model ${\cal M}$ is the minimal length—across all possible formulations (in any language) of ${\cal M}$—of all the assumptions of ${\cal M}$ that are needed to derive all the available results of ${\cal M}$. The assumptions must contain all references needed to ensure that any comparison between the empirical data $D$ and the corresponding results of ${\cal M}$ are measurable [^35] For a discussion of the concept of “measurable” see App. A and (Scorzato, 2013).. The (epistemic) simplicity of a model ${\cal M}$ is the inverse of its complexity.
Being defined as the minimum across all possible formulations, Def. 1 is obviously language independent. It is a property of the model and it therefore depends on the concepts that define the empirical content of the model, but it does not have any catastrophic dependence on the language. In fact, the interesting observation is that it is also in non-trivial, in general, thanks to the analysis in App. A.
@@ Appendix C Alternative notions of complexity
Kolmogorov complexity is a very popular measure of complexity, also widely used to estimate the complexity of a theory (Chaitin, 1975). The analysis in (Zenil, 2017) goes into great details to justify why Kolmogorov complexity is well suited to estimate the complexity of both scientific and any ML model. Nevertheless, it is not the only possibility and one may question how the present analysis would change if we had used a different measure. For example, any specific compression software, like, e.g., gzip, implicitly defines an alternative notion of complexity. However, in all these cases, our conclusions about the complexity of DNN models would not change [^40] See, however the discussion in (Zenil, 2020b) that emphasises how compression techniques miss important features that are captured by Kolmogorov complexity.. Counting the atomic sentences (Carey, 2023) would also not change our conclusions.